Anne and Nancy use a metal alloy that is 15.6% copper to make jewelry. How many ounces of an alloy that is 14% copper must be mixed with an alloy that is 16% copper to form 100 ounces of the desired alloy?
Answer by mananth(16946)

percent ---------------- quantity
Metal alloy I 14 ---------------- x ounces
Metal alloy II 16 ------ 100 - x ounces
Mixture 16 ---------------- 100

14x+16(100-x)=15.6*100
1400x+1600(100- x)=156000
1400x+160000-1600x=156000
1400x-1600x= 156000 -160000
-200 x = -4000
/ -200
x=20ounces 14.00% Metal alloy I
80 ounces 16.00% Metal alloy II

m.ananth@hotmail.ca